What is the pressure in kilopascals in a 35.0 L balloon at 25 °C filled with pure hydrogen
gas produced by the reaction of 34.11 g of CaH2 with water?
1
Expert's answer
2019-06-21T08:46:17-0400

Solution: CaH2 + 2H2O = Ca(OH)2 + H2

From one mol of CaH2 you get 1 mol H2. The amount of CaH2 is 34.11/42.08=0.81 mol, and it equals to the amount of hydrogen. 25 °C is 25+273.15=298.15 °K. The preasure in the baloone can be calculated by the ideal gas law: P=nRT/V, so P=0.81*8.314*298.15/35.0=57.37 kPa.

Answer: 57.37 kPa.